Tom and Jerry together had 115. After tom had spent 10, his money was 3/4 that of jerry. How much did Jerry have? (Show your solution)
Answer by mananth(16949)

Tom and Jerry together had 115. After tom had spent 10, his money was 3/4 that of jerry. How much did Jerry have? (Show your solution)
..
let jerry have x
then tom has 115-x
..
tom spends 10
so he has 115-x-10
=105-x
..
105-x= 3/4 * x
multiply by 4
4(105-x)=3x
420-4x=3x
420 = 7x
/7
60=x
...
Jerry had 60
...
m.ananth@hotmail.ca